OpenStudy (anonymous):

solve the equation for y. 7x-3y=4

OpenStudy (anonymous):

7x-3y=4 you cant solve this, but you can isolate for x and y (i assume for y). -3y=4-7x 3y=7x-4 y=7/3x-4/3 and for x 7x-3y=4 7x=3y+4 x=3/7y+4/7